Equation 1: y= 4x-11
Equation 2: 2x+y=7
What is answer solved by two tables, graphing and subsitution

Answer by rfer(16322)

2x+4x-11=7
6x=18
x=3
-------------
y=4(3)-11
y=1
--------------
proof
1=4(3)-11